OpenStudy (anonymous):

Covert each angle to decimal degrees. Round answers to four significant digits. 2) 42degrees10' 4) 81degrees52'30" 6) 35degrees15' 8) 24degrees0'45' 10) 60degrees40'40"

OpenStudy (amilapsn):

hint: 1'=1/60 degrees 1''=(1/60)'

OpenStudy (lynfran):

divide the minutes by 60 divide the seconds by 3600

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

So the first one i got 42degrees 1/10

OpenStudy (lynfran):

i will do the last one